Write the slope-intercept form of the equation of the line passing through point (-4,0) and parallel to the line y=3x+2
a. y=3x+12 b. y=-3x+4 c. y=-3x-12 d. y=3x-4

Answer by stanbon(75887)

Write the slope-intercept form of the equation of the line passing through point (-4,0) and parallel to the line y=3x+2
-------
slope = 3
Form: y = mx + b
0 = 3*(-4) + b
b = 12
-----
Equation:
y = 3x +12
=====================
Cheers,
Stan H.
=====================

a. y=3x+12 b. y=-3x+4 c. y=-3x-12 d. y=3x-4